OpenStudy (anonymous):

Mitchell deposited $651.37 in a savings account that earns 3.9% simple interest. What is Mitchell’s account balance after 15 years?

OpenStudy (anonymous):

I need help on this i honestly don't even understand anything that is happening in this class please help me!(O.O)

OpenStudy (campbell_st):

use the formula I = Prn P = principal ( amount invested) $651.37 r = interest rate as a decimal 3.9% = 0.039 n = number of time periods 15 just substitute andevaluate for an answer.

OpenStudy (anonymous):

Thank you so much =D i got 381

OpenStudy (campbell_st):

I think you may need to include the 5 cents in the answer $381.05 this is the interest the account balance will be the Principal + Interest

OpenStudy (anonymous):

yeah i did=D Thanks

OpenStudy (anonymous):

oh it this the same way Stacey currently has an account balance of $1,007.78. Her initial deposit on the account was $616 and it earned 5.3% simple interest. How long has Stacey held the account?

OpenStudy (campbell_st):

so in this question you need to find the interest 1007.78 - 616 = 391.78 when you know the interest I you need to formula I = Prn or 391.78 = 616 * 0.053* n you need to solve for n

OpenStudy (anonymous):

31 years?=D

OpenStudy (campbell_st):

ummm the question becomes \[\frac{391.78}{(616 \times 0.053)} = \]

OpenStudy (anonymous):

oh ok 12 years!=D